Hi this is a system of equations with 3 variables.
4x-3y-2z=-5
3x+2y+7z=9
-2x+5y+8z=1
What are the values for x, y, and z? Thank you.

Answer by Alan3354(69443)

4x-3y-2z=-5
3x+2y+7z=9
-2x+5y+8z=1
-----------
I'll use determinants, since you didn't mention a method.
--------------
Det = 4(16-35) + 3(24+14) - 2(15+4)
= -76 + 114 - 38
Det = 0
No unique solution.